Marydid not like to tell her age so her mother answered for her. I am just 7 times as old as she is now. In 20 years she will be half the age that I will be then. How old is Mary?
Answer by w_parminder(53)

Let Mary's present age = X years
So her mother's present age = 7X
After 20 years,
Mary's age = X + 20
Mother's age = 7X + 20
According to question (A.T.Q)
after 20 years, Mary will be half the age of her mother, OR
her mother's age will be the double of Mary's age
i.e. 7X + 20 = 2(X + 20)
7X + 20 = 2X + 40
5X = 20
So, X, i.e. Mary's present age, = 20/5 = 4 years